OpenStudy (anonymous):

for parallelogram LMNO, if m angle L =5x-28 and m angle N = 3x-10,find m angle O

OpenStudy (anonymous):

angleL = angleN as opposite angles in a parallelogram are equal so 5x-28 = 3x-10 gives 2x = 18 x = 9 so angleL = 5(9) -28 = 45 -28 = 17 Now angleL and angleO are adjacent angles and in a ||gm adjacent angles total to 180 degree so angleO = 180 - angleL = 180 - 17 = 163 deg

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thx !!!